OpenStudy (anonymous):

a man wishes to enclose 2 separate lots with 300 yrds of fencing, one lot is a square n the other a rectangle twice as long as its wide. determine the dimensions of each lot if total area to be enclosed is a minimum. use the quadratic relation to hsow ur work

OpenStudy (anonymous):

a fig. would help!

OpenStudy (anonymous):

iam sorry?!!

OpenStudy (anonymous):

do u know the formule!

OpenStudy (anonymous):

plz if u r not going to answer then dont bother wasting my time n urs. THNK U

OpenStudy (anonymous):

@Rohangrr UR TECHNICALLY A FIG!

OpenStudy (anonymous):

@khadija UR jack of all trade but master of none! seriously u r the biggest joker out here! lol! ha hah ha

OpenStudy (anonymous):

@Rohangrr u've proved me right! u r an idiot of life! this is wat u call NO LIFE! am sorry:p jokes on u!

OpenStudy (anonymous):

dont flirt the BIGGEST JOKER OS!! Sorry AH! sorry:p jokes on JOKER @khadija

OpenStudy (anonymous):

Hello guys, Please be respectful.

OpenStudy (anonymous):

OK tht was just simply stupid!! n plz if u dont wish to answer u can leave no1 asked u!! those who want to will help nn a fool like u wud waste time

OpenStudy (anonymous):

hey in this case let us assume length of one side of square be x therefore total fencing for square would be 4x while a rectangle having its length double of of its width hey @khadija is there any relation between square and rectangle

OpenStudy (anonymous):

@sheg NOPE:(

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i feel it would be 300 yards each

OpenStudy (anonymous):

nope..

OpenStudy (anonymous):

or width should be equal to side of the square

OpenStudy (anonymous):

??? u r finding the dimensions

OpenStudy (kropot72):

I have a solution: The square has sides of 35.29 yards The rectangle has short sides 26.47 yards and long sides 52.94 yards. Do you want me to show my workings?

OpenStudy (anonymous):

@kropot72 yes please, tht wud be great!! n thnk u=D ur help is greatly appreciated

OpenStudy (anonymous):

@kropot72 can u help me with 2 moreas ell. If u can =D thnnkk u

OpenStudy (kropot72):

Let length of side of square = s Let length of short side of rectangle = r 4s + 2r + 4r = 300 4s = 300 - 6r s = 75 -1.5r Area of square = r^2 = (75 - 1.5r)^2 Area of rectangle = 2r^2 Let total area = A A = (75 - 1.5r)^2 + 2r^2 = 4.25r^2 - 225r + 5625 Differentiating A w.r.t. r: dA/dr = 8.5r - 225 Put the result of differentiating equal to zero and solve to find the minimumvalue of A: 8.5r - 225 = 0 8.5r = 225 r = 26.47 yards Substituting for r gives: s = 75 - (26.47 * 1.5) = 35.29 yards The square has sides of 35.29 yards The rectangle has short sides 26.47 yards and long sides 52.94 yards.

OpenStudy (anonymous):

ohh nice!! THNNNKKKK UUUUU SO MUCHHH!! tht was really helpfull :D thnx;)

OpenStudy (kropot72):

You're welcome. Sorry I can't help with more. Must get some sleep!!!

OpenStudy (anonymous):

aww no its ok dw:) ths was awesome =D go get ur rest!!

OpenStudy (farmdawgnation):

Rohangrr, do not call other users jokers and insult them. It's against our Code of Conduct. Thanks!